OpenStudy (anonymous):

what is the solution to the system? 2x+5y=-4 4x-5y=22 the options are (3,-2) (13/3, -25/3), (-2,3) (3, 2/5)

OpenStudy (anonymous):

2x + 5y = -4 4x - 5y = 22 ---------------add 6x = 18 x = 3 now sub 3 in for x in either one of the equations above 2x + 5y = -4 2(3) + 5y = -4 6 + 5y = -4 5y = -4 - 6 5y = -10 y = -2 answer : (3,-2) any questions ?

OpenStudy (anonymous):

no, thank you so much! im gonna use this to teach myself on another problem that i have to solve the same way. thank you!